# Universality of high-dimensional scaling limits of stochastic gradient descent

### v1

@@ 4.1. Ballistic $\delta$-localizability on the coordinate-delocalized set
We start by using the closeness of expectations from Section 3 to show that the same strong $\delta$-localizability moment conditions that ensure ballistic limits for Gaussian projective models also hold for the $\nu$-MM data as long as the parameter $\Theta$ is coordinate-delocalized. We note that without the restriction to coordinate-delocalized parameter, the inner product of the gradient with certain directions may blow up in the limit.
@@
Item 2. The delicate term to handle in this way is the first part of item 2. For instance, for $\Theta\in{\mathscr{D}}_{\zeta}$, let us consider
@@ 4.2. The parameter stays coordinate-delocalized for linear timescales
Let $\tau_{R}$ be as in (4.1), and for $\zeta<\zeta_{0}\leq 1/8$, let $\tau_{\zeta_{0}}=\min\{\ell:\Theta\notin{\mathscr{D}}_{\zeta_{0}}\}$, and $\tau=\min\{T\delta^{-1},\tau_{R},\tau_{\zeta_{0}}\}$. Consider the evolution of the $\ell^{3}$ norm of the parameter. For any $i\in[d]$,
@@
where $M_{\ell}^{a}$ is a martingale. Taking absolute values of both sides, and then using Lemma 4.3 to replace the expectation above up to an $O(d^{-1+4\zeta_{0}})$ error, for $\ell\leq\tau$,
@@
We bound the martingale term by applying Doob’s maximal inequality, the Burkholder–Davis–Gundy inequality, and Minkowski’s integral inequality in turn to find that for each $a\in[k_{1}]$ and $p\geq 1$ we have
@@
For any $p\leq 10$, the first term on the right-hand-side is $O(1)$ uniformly over $\Theta\in{\mathscr{D}}_{\zeta_{0}}\cap B_{R}(0)$ by Corollary 3.3; and since we assume at least $20$ moments on $\nu$, the second term is also seen to be $O(1)$. Plugging in $r=\epsilon d^{-1/2+\zeta_{0}}$ for $\epsilon$ sufficiently small (depending on the constants $R,T$, etc.) and using that $\delta\asymp 1/d$, for every $p\leq 10$,
$\mathbb{P}(\max_{\ell\leq\tau}|M_{\ell}^{a}|>\epsilon d^{-1/2+\zeta_{0}})\leq O(r^{-p}\delta^{p}d^{p/2})=O(d^{-10\zeta_{0}})\,.$
Using a union bound over $i\in[d]$ and $a\in[k_{1}]$, we get that with probability $1-O(d^{-10\zeta_{0}+1})$, which is 1o(1)1-o(1) if $\zeta_{0}>1/10$, the following holds for all initializations in ${\mathscr{D}}_{\zeta}\cap B_{R}(0)$ for $d$ sufficiently large: For all time steps $\ell\leq\tau$, and all $i\in[d]$,
$\max_{a}|m_{ai}(\ell)|\leq 2\epsilon d^{-1/2+\zeta_{0}}+C\delta\sum_{k\leq\ell}\max_{a}|m_{ai}(k)|\,$,
where the extra $\epsilon d^{-1/2+\zeta_{0}}$ was used to absorb both $\max_{a}|m_{ai}(0)|$, since $\zeta_{0}>\zeta$, and the $O(d^{-1+4\zeta_{0}})$ error since $\zeta_{0}\leq 1/8$. Then by the discrete Gronwall inequality, we deduce that for $\zeta<\zeta_{0}$ with $\zeta_{0}\in(1/10,1/8]$, with probability 1o(1)1-o(1), if the initialization and $\bm{\mu}$ are in ${\mathscr{D}}_{\zeta}$, then for all $i\in[d]$, and all $a\in[k_{1}]$
$\sup_{\ell\leq\tau}|\theta^{a}_{i}(\ell)|\leq 2\epsilon d^{-1/2+\zeta_{0}}e^{C^{\prime}T}\,.$
Taking any $\zeta_{0}\in(\zeta\vee\frac{1}{10},\frac{1}{8}]$, for fixed $T$, taking $\epsilon$ sufficiently small, this is bounded by $d^{-1/2+\zeta_{0}}$. ∎
@@ 5. Universality of ballistic dynamics
We now show that the martingale term is negligible. To this end, first note that for the first term, the same bound in [9, Eq. 6.4] applies. In fact, due to part 1 of item 3 of strong $\delta$-localizability (which holds for all $\ell\leq\tau\wedge T\delta^{-1}$ by Lemma 4.2), we can improve that bound to
@@
Next we deal with each part of the error term, $\mathcal{E}_{\ell}$, in turn. For the first part of the error term, we may apply [9, Eq. 6.5] directly. For the second part, by strong localizability, we can improve [9, Eq. 6.6] as well to get
@@ 6. Non-universality of diffusive limits
Since there are many terms appearing in the difference in expectations, in order to not have to worry about possible cancellations between them, we will expand in orders of $\rho$ and by taking $\rho$ to be a small constant, establish a distinction.
@@
Suppose that the pair of summary statistics $\widetilde{\mathbf{G}}=(\sqrt{d}\langle\theta,\theta_{*}\rangle,\sqrt{d}(\|\theta\|_{2}^{2}-R_{*}))$ have an SDE limit on linear timescales, as otherwise the claim that the limit is not the Gaussian one holds vacuously. Consider the summary statistic $\tilde{u}(\theta)=\sqrt{d}\langle\theta,\theta_{*}\rangle$. If $\tilde{u}_{\ell}=\tilde{u}(\theta_{\ell})$, then

### v2

@@ 4.1. Ballistic $\delta$-localizability on the coordinate-delocalized set
We start by using the results of Section 3 to show that strong $\delta$-localizability also holds for the $\nu$-MM data when $\Theta$ is coordinate-delocalized.
@@
Item 2. The delicate term to handle is the first part of item 2. For instance, for $\Theta\in{\mathscr{D}}_{\zeta}$, consider
@@ 4.2. The parameter stays coordinate-delocalized for linear timescales
Let $\tau_{R}$ be as in (4.1), and for $\zeta<\zeta_{0}\leq 1/8$, let $\tau_{\zeta_{0}}=\min\{\ell:\Theta\notin{\mathscr{D}}_{\zeta_{0}}\}$, and $\tau=\min\{T\delta^{-1},\tau_{R},\tau_{\zeta_{0}}\}$. Consider the evolution of the parameter’s $\ell^{3}$ norm. For any $i\in[d]$,
@@
where $M_{\ell}^{a}$ is a martingale. Take absolute values of both sides. Then using Lemma 4.3 to replace the expectation above up to an $O(d^{-1+4\zeta_{0}})$ error, for $\ell\leq\tau$,
@@
Bound the martingale term by applying Doob’s maximal inequality, the Burkholder–Davis–Gundy inequality, and Minkowski’s integral inequality in turn to find that for each $a\in[k_{1}]$ and $p\geq 1$,
@@
For any $p\leq 10$, the first term on the right-hand-side is $O(1)$ uniformly over $\Theta\in{\mathscr{D}}_{\zeta_{0}}\cap B_{R}(0)$ by Corollary 3.3; and since we assume at least $20$ moments on $\nu$, the second term is also seen to be $O(1)$. Plugging in $r=d^{-\xi}d^{-1/2+\zeta_{0}}$ for $\xi$ to be determined sufficiently small, and using that $\delta=O(1/d)$ and $T=O(\log d)$, for $p=10$,
(4.2) $\mathbb{P}(\max_{\ell\leq\tau}|M_{\ell}^{a}|>d^{-1/2+\zeta_{0}-\xi})\leq O(r^{-p}\delta^{p/2}(\log d)^{p/2})=O(d^{10\xi}d^{-10\zeta_{0}}(\log d)^{5})\,.$
Using a union bound over $i\in[d]$ and $a\in[k_{1}]$, we get that with probability $1-\tilde{O}(d^{10\xi}d^{-10\zeta_{0}+1})$—which is 1o(1)1-o(1) if $\zeta_{0}>1/10$ and $\xi$ is sufficiently small—the following holds for all initializations in ${\mathscr{D}}_{\zeta}\cap B_{R}(0)$ for $d$ sufficiently large: For all time steps $\ell\leq\tau$, and all $i\in[d]$,
$\max_{a}|m_{ai}(\ell)|\leq 2d^{-1/2+\zeta_{0}-\xi}+C\delta\sum_{k\leq\ell}\max_{a}|m_{ai}(k)|\,.$
Here, the extra $d^{-1/2+\zeta_{0}-\xi}$ was used to absorb both $\max_{a}|m_{ai}(0)|$ and the $O(d^{-1+4\zeta_{0}})$ error since $\zeta<\zeta_{0}\leq 1/8$ and $\xi$ is sufficiently small. Then by the discrete Gronwall inequality, we deduce that for $\zeta<\zeta_{0}$ with $\zeta_{0}\in(1/10,1/8]$, with probability 1o(1)1-o(1), if the initialization and $\bm{\mu}$ are in ${\mathscr{D}}_{\zeta}$, then for all $i\in[d]$, and all $a\in[k_{1}]$
(4.3) $\sup_{\ell\leq\tau}|\theta^{a}_{i}(\ell)|\leq 2d^{-1/2+\zeta_{0}-\xi}e^{C^{\prime}T}\,.$
Taking any $\zeta_{0}\in(\zeta\vee\frac{1}{10},\frac{1}{8}]$, for $T=\varepsilon\log d$, with $\varepsilon$ sufficiently small, this is bounded by $d^{-1/2+\zeta_{0}}$. ∎
Proof of Theorem 1.4.
By Lemma 4.1, with probability 1o(1)1-o(1), one has $\Theta_{\ell},\bar{\Theta}_{\ell}\in\mathcal{D}_{1/8}$ for all $\ell\leq\varepsilon d\log d\wedge\tau_{R}\wedge\bar{\tau}_{R}$. This uniformly bounds differences of expectations for functions of $\bm{\theta}^{\top}X$ with $\|f\|_{\mathcal{F}_{0}}\leq 1$ by $\|\bm{\theta}\|_{3}^{3}\leq d^{-1/2+3\zeta}=o(1)$ by Lemma 3.1 under Assumption 1. This concludes the proof because $\|f\|_{C^{3}_{b}}=\|f\|_{\mathcal{F}_{0}}$. ∎
@@ 5. Universality of ballistic dynamics
We now show that the martingale term is negligible. To this end, first note that for the first term, the same bound in [10, Eq. 6.4] applies. In fact, due to part 1 of item 3 of strong $\delta$-localizability (which holds for all $\ell\leq\tau\wedge T\delta^{-1}$ by Lemma 4.2), we can improve that bound to
@@
Next we deal with each part of the error term, $\mathcal{E}_{\ell}$, in turn. For the first part of the error term, we may apply [10, Eq. 6.5] directly. For the second part, by strong localizability, we can improve [10, Eq. 6.6] as well to get
@@ 6. Non-universality of diffusive limits
Since there are many terms appearing in the difference in expectations, in order to not have to worry about possible cancellations between them, we will expand in orders of $\rho$ and establish a distinction by taking $\rho$ small enough.
@@
Suppose that the pair of summary statistics $\widetilde{\mathbf{G}}=(\sqrt{d}\langle\theta,\theta_{*}\rangle,\sqrt{d}(\|\theta\|_{2}^{2}-R_{*}))$ have an SDE limit on linear timescales, as otherwise the claim that the limit is not the Gaussian one holds vacuously. Consider the summary statistic $\tilde{u}(\theta)=\sqrt{d}\langle\theta,\theta_{*}\rangle$. If $\tilde{u}_{\ell}=\tilde{u}(\theta_{\ell})$,